# TimeDistill: Efficient Long-Term Time Series Forecasting with MLP via Cross-Architecture Distillation

### v1

@@ 6 Conclusion and Future work
We propose TimeDistill, a cross-architecture KD framework enabling lightweight MLP to surpass complex teachers. By distilling multi-scale and multi-period patterns, TimeDistill efficiently transfers temporal and frequency-domain knowledge. Theoretical interpretations and experiments confirm its effectiveness. Future work includes distilling from advanced time series models, e.g. time series foundation models, and incorporating multivariate patterns.
7 Impact Statement
Transparency Impact. Transparency is vital to facilitate accountability, competition, and collective understanding (Bommasani et al., 2024). To support these objectives, we will publicly release our code base, data sources, and evaluation pipeline. However, due to the requirements of anonymized submission, these materials are currently included only in the supplementary material.
Environmental Impact. TimeDistill achieves fast inference speed, which significantly reduces its carbon footprint by lowering GPU usage. This reduction also contributes to decreased carbon emissions from the local electricity grid.
Ethical Impact. While TimeDistill performs well in long-term time series forecasting, caution is needed in critical and high-stakes domains like healthcare, where incorrect predictions could have severe consequences.
Appendix A Related Work
A.1 Debate in Long-Term Time Series Forecasting
The trade-off between performance and efficiency has prompted a long-standing debate between Transformer-based models and MLP in long-term time series forecasting (Zeng et al., 2023; Zhang et al., 2022; Lin et al., 2024). Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022) were among the leading Transformer-based methods. However, recent findings show that a simple Linear or MLP model can achieve performance comparable to these complex Transformer models across various benchmarks while offering significantly better efficiency (Zeng et al., 2023). This outcome has raised questions about the necessity of Transformers in time series forecasting. Following this, research has moved in two directions. One direction suggests that the issues with Transformer-based models arise from the way they are applied. For example, PatchTST (Nie et al., 2023) uses patching to preserve local information, and iTransformer (Liu et al., 2024) focuses on capturing multivariate correlations. These approaches surpass the simple one-layer MLP and demonstrate that Transformers can still deliver strong results in time series forecasting if applied effectively. Meanwhile, CNN-based models have also shown strong performance similar to Transformer-based models. TimesNet (Wu et al., 2022) transforms 1D time series into 2D variations and applies 2D CNN kernels, MICN (Wang et al., 2023) adopt multi-scale convolution structures to capture local features and global correlations, and ModernTCN (Luo & Wang, 2024) proposes a framework with much larger receptive fields than prior CNN-based structures. Nevertheless, these powerful CNN-based models also face efficiency issues, which further broadens the scope of the debate between performance and efficiency. The other direction focuses on developing lightweight MLP, such as N-BEATS (Oreshkin et al., 2020), N-hits (Challu et al., 2023), LightTS (Zhang et al., 2022), FreTS (Yi et al., 2023), TSMixer (Chen et al., 2023), TiDE (Das et al., 2023), which offer improved efficiency. However, these models typically only match, rather than surpass, the performance of state-of-the-art Transformer-based methods and CNN-based methods. In summary, while Transformer-based and CNN-based models generally offer better performance, simple MLP is more efficient. Therefore, we managed to combine the performance of Transformer-based and CNN-based models with MLP to produce a powerful and efficient model by cross-architecture KD.
A.2 Knolwedge Distillation on Time Series
Knowledge distillation (KD(Hinton, 2015) transfers knowledge from a larger, more complex model (teacher) to a smaller, simpler model (student) while maintaining comparable performance. By aligning the output distributions of teacher and student models, KD provides richer training signals than hard labels alone, enabling the student to capture subtle patterns that the teacher has learned. In the context of time series, CAKD (Xu et al., 2022) introduces a two-stage distillation scheme that distills features using adversarial and contrastive learning and performs prediction-level distillation. LightTS (Campos et al., 2023) designs a KD framework specifically for cases where the teacher is an ensemble classifier in time series classification tasks, which limits its applicability to teachers with other architectures. Both of these works do not incorporate time series-specific designs. In contrast, our framework emphasizes extracting essential time series patterns, including multi-scale and multi-period patterns, enabling more effective knowledge distillation. Additionally, we address architectural differences and are the first to explore cross-architecture KD.
Appendix B Theoretical Understanding from Data Augmentation Perspective
In this section, we propose a theoretical viewpoint that TimeDistill unifies the idea of augmenting training data with blended data that combines teacher outputs and ground truth. This concept is closely related to mixup (Zhang, 2017), a well-known data augmentation strategy. Mixup creates new training samples by mixing two examples’ inputs and corresponding targets according to a random interpolation coefficient. Analogously, in multi-scale or multi-period distillation, we create augmented samples that blend teacher predictions and ground truth signals in a manner that provides richer supervision signals. Concretely, in a standard mixup procedure, one obtains augmented samples $(\tilde{x},\tilde{y})$ from two distinct training samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$ by drawing a mixing ratio $\lambda$ and computing:
@@ Appendix B Theoretical Understanding from Data Augmentation Perspective
$\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug}$ (13)
when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (Zhang, 2017) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@
where $x_{0}=Y$ and $x_{m+1}=\hat{Y}_{t}^{(m)}$ for $m=0,\dots,M$. By Jensen’s inequality (Jensen, 1906), for a convex function $f$, we have:
@@
$\mathcal{L}_{sup}+\eta\mathcal{L}_{period}\geq\mathcal{L}_{aug}$ (14)
@@
When the temperature $\tau$ goes to $\infty$, minimizing KL divergence between two distributions is equivalent to minimizing the MSE error (Kim et al., 2021b). Thus, we have
@@
and the log sum inequality (Thomas & Joy, 2006) states that
@@
Table 6: Dataset characteristics for time series long-term forecasting task.
@@
*The forecastability is borrowed from TimeMixer (Wang et al., 2024a). A larger value indicates better predictability.
Appendix C Implementation Details
Datasets Details. We conduct experiments to evaluate the performance and efficiency of TimeDistill on eight widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following the pipeline in (Liu et al., 2024; Wang et al., 2024a; Luo & Wang, 2024). Detailed descriptions of these datasets are provided in Table 6.
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following (Liu et al., 2024; Luo & Wang, 2024; Wang et al., 2024a, 2023; Wu et al., 2022, 2021; Zhou et al., 2021; Zeng et al., 2023):
$\text{MSE}=\frac{1}{S\times C}\sum_{i=1}^{S}\sum_{j=1}^{C}(Y_{ij}-\hat{Y}_{ij})^{2},$ (15)
$\text{MAE}=\frac{1}{S\times C}\sum_{i=1}^{S}\sum_{j=1}^{C}|Y_{ij}-\hat{Y}_{ij}|.$ (16)
@@ Appendix C Implementation Details
Table 7: Experiment configuration of TimeDistill.
@@
Experiment Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on a cluster equipped with NVIDIA A100, V100, and K80 GPUs. The teacher models are trained using their default configurations as reported in their respective papers. For the student MLP model, we employ a combination of a decomposition scheme (Zeng et al., 2023; Wu et al., 2021; Zhou et al., 2022) and a two-layer MLP with hidden dimension $D$ as the architecture. When using TimeDistill for distillation, the teacher model is frozen, and only the student MLP is trained. The initial learning rate is set to 0.01, and the ADAM optimizer (Kingma, 2014) is used with MSE loss for the training of the student model. We apply early stopping with a patience value of 5 epochs. The batch size is set to 32. The temperature for multi-period distillation is set to $\tau=0.5$. The number of scales is set to $M=3$. We use two types of normalization methods, i.e. non-stationary (Liu et al., 2022) and revin (Kim et al., 2021a). We perform a hyperparameter search for $\alpha$ and $\beta$ within the range {0.1, 0.5, 1, 2}. We conduct hyperparameter sensitivity analysis in Appendix G. For consistency, TimeDistill defaults to using ModernTCN as the teacher and fixes the look-back window length to 720. Additional experiments with other teacher models and different look-back window lengths are presented in Section 5.3. Additional detailed model configuration information is presented in Table 7.
Channel Independent MLP. We denote the model as $f$. The channel-independent strategy predicts each channel independently, defined as:$\hat{Y}=[f(x^{1}),\dots,f(x^{C})]$, where $\hat{Y}=[\hat{y}^{1},\dots,\hat{y}^{C}]$ denotes the model prediction. In contrast, the channel-dependent strategy considers the combination of all channels simultaneously, defined as: $\hat{Y}=f(x^{1},\dots,x^{C})$. In this paper, for MLP, we adopt the channel-independent strategy because it eliminates the need to model inter-channel relationships (Han et al., 2024), which reduces model complexity and results in a more lightweight model, which is especially advantageous when dealing with a large number of variables.
Appendix D Implementation Details for Preliminary Study.
@@
$\mathbf{\hat{Y}}^{m}=\mathrm{Conv}(\mathbf{\hat{Y}}^{m-1},\mathrm{stride}=2),$ (17)
@@
Table 8: Performance promotion obtained by our TimeDistill framework with different teacher models. $\Delta_{MLP}$ ($\Delta_{Teacher}$) represents the performance promotion between MLP+TimeDistill and a trained MLP (Teacher). We report the average performance from all prediction lengths.
@@
(2024) (2024) (2024a) (2023) (2023) (2022) (2022) (2021)
@@
Appendix E Different Teachers Analysis
@@
On Traffic dataset, which involves a large number of channels (861 channels) and is particularly challenging for channel-independent models like the student MLP, TimeDistill achieves the second-best performance, improving MSE by 10.30% when the teacher is iTransformer. This is likely due to TimeDistill can implicitly distill iTransformer’s knowledge of modeling multivariate correlations to student MLP. We further explore whether TimeDistill can implicitly learn multi-variate correlation via distillation in Appendix J.
@@
Appendix F Different Students Analysis
@@
Appendix G Hyperparameter Sensitivity Analysis
@@
Figure 9: Sensitivity Analysis Result of Predcition Level $\alpha$ and Feature Level $\beta$ Distillation on Different Datasets.
@@
Figure 10: Hyperparameter sensitivity with respect to the number of scales $M$ and temperature $\tau$ on ECL dataset. The results are recorded with the lookback window length T = 720 and averaged across all prediction window lengths $S~{}\in\{96,192,336,720\}$.
Table 9: Results of directly match the patterns in ground truth time series on eight datasets. The look-back length is set to be consistent at 720 for fair comparison. The MSE and MAE metrics are averaged from all prediction lengths $S\in\{96,192,336,720\}$.
@@
Table 10: Static and runtime metrics of TimeDistill, MLP, and other mainstream models on the ECL dataset. For fair comparison, the look-back length for each model is set to be consistent at 720 and the batch size is set to be 16. The static and runtime metrics are averaged from all prediction lengths $S~{}\in\{96,192,336,720\}$.
@@
iTransformer (Liu et al., 2024) 0.163 0.258 3.6377 5,276,496
ModernTCN (Luo & Wang, 2024) 0.167 0.262 6.2233 132,075,892
TimeMixer (Wang et al., 2024a) 0.165 0.259 7.7662 5,064,993
PatchTST (Nie et al., 2023) 0.165 0.266 2.9802 24,949,584
MICN (Wang et al., 2023) 0.181 0.293 2.715 60,215,726
Crossformer (Zhang & Yan, 2023) 0.203 0.299 8.7768 2,264,036
Fedformer (Zhou et al., 2022) 0.274 0.376 32.2642 16,827,399
Autoformer (Wu et al., 2021) 0.238 0.347 196.5266 14,914,398
@@
Appendix H Directly Match the Patterns in Ground Truth Time Series
@@
$\mathcal{L}^{\prime}_{sup}=\mathcal{L}_{scale}^{sup}+\mathcal{L}_{period}^{sup}.$ (18)
@@
Appendix I Efficiency Analysis
In the main text, we have presented the efficiency analysis in Figure 2. Here, we provide the quantitative results in Table 10. Latency-sensitive applications require fast inference, making it crucial for models to minimize inference time. However, most works in the time series domain primarily focus on less time-sensitive training time comparisons (Liu et al., 2024; Lin et al., 2024). Additionally, edge-device applications demand models with fewer parameters for deployment. Therefore, in our experiments, we emphasize comparing both inference time and model parameters to evaluate practical efficiency. For a fair comparison, we fix the batch size to 16 during inference and record the average inference time per batch. Notably, TimeDistill (MLP) demonstrates a significant efficiency advantage over Transformer-based models, such as iTransformer, PatchTST, Crossformer, FEDformer, and Autoformer, as well as CNN-based models like ModernTCN and MICN, and the complex MLP-based model TimeMixer, which employs multiple MLPs. In terms of inference time, TimeDistill achieves up to 196$\times$ faster inference than other methods. Additionally, it requires up to 60$\times$ fewer parameters compared to other methods. Beyond the ECL dataset, we also visualize efficiency analyses for other datasets in Figure 14. Across eight datasets, TimeDistill consistently resides in the lower-left corner, achieving the best trade-off between efficiency (inference time) and performance (MSE).
Figure 11: Multi-Variate Person Correlation Visualization on ETTh1 and ETTh2 datasets.
Appendix J Can TimeDistill Learn Multi-Variate Correlation via Distillation?
Some teacher models, such as iTransformer and ModerTCN, are designed in a channel-dependent (CD) manner. While this approach can utilize multi-variate correlation effectively, it incurs a higher computational burden. In contrast, advanced methods like PatchTST, TimeMixer, and our method TimeDistill are employed in a channel-independent (CI) manner, which offers better computational efficiency. This leads to an intriguing question: Can TimeDistill enable MLP to learn multi-variate correlation from the teacher without an explicit distillation design? To explore this, we present a preliminary study using Figure 11, which illustrates two cases from the ETTh1 and ETTh2 datasets, each containing seven variables. Our observations reveal that TimeDistill’s multi-variate correlation more closely resembles that of the teacher compared to the MLP model before distillation. This suggests that TimeDistill facilitates learning underlying structures among multi-variates through distillation even though the student MLP is a CI model. These findings highlight a potential benefit of distillation that merits further exploration. We leave this for further work.
@@ Appendix I Efficiency Analysis
Table 11: Performance promotion obtained by our TimeDistill framework with different student models. The notation nL-d represents the structure of an MLP model, where nL specifies the number of layers and d denotes the dimension of the hidden units in each layer. For instance, 2L-512 refers to a model with 2 layers and a hidden dimension of 512. We report the average performance from all prediction lengths. $\Delta_{Student}$ represents the performance promotion between Student+TimeDistill and original trained student.
Student Models MLP MLP MLP MLP LightTS TSMixer
(2L-512) (3L-512) (4L-512) (2L-1024) (2022) (2023)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 Student 0.502 0.489 0.487 0.497 0.468 0.481 0.495 0.500 0.465 0.471 0.471 0.474
+TimeDistill 0.428 0.445 0.442 0.448 0.443 0.453 0.442 0.446 0.436 0.445 0.433 0.446
$\Delta_{Student}$ 14.74% 9.00% 9.11% 9.86% 5.29% 5.70% 10.79% 10.83% 6.26% 5.57% 8.02% 5.92%
ETTh2 Student 0.393 0.438 0.742 0.601 0.908 0.668 0.639 0.563 0.675 0.582 0.376 0.420
+TimeDistill 0.345 0.397 0.341 0.390 0.362 0.406 0.344 0.390 0.382 0.423 0.351 0.399
$\Delta_{Student}$ 12.16% 9.46% 53.97% 35.05% 60.12% 39.19% 46.20% 30.72% 43.42% 27.26% 6.83% 5.01%
ETTm1 Student 0.391 0.413 0.378 0.406 0.392 0.414 0.377 0.402 0.376 0.399 0.370 0.395
+TimeDistill 0.354 0.390 0.346 0.380 0.353 0.389 0.346 0.380 0.358 0.382 0.354 0.383
$\Delta_{Student}$ 9.42% 5.65% 8.54% 6.34% 9.90% 6.12% 8.40% 5.56% 4.83% 4.29% 4.13% 2.92%
ETTm2 Student 0.300 0.373 0.359 0.408 0.346 0.393 0.307 0.374 0.283 0.346 0.295 0.347
+TimeDistill 0.252 0.316 0.259 0.321 0.254 0.317 0.245 0.312 0.258 0.324 0.253 0.319
$\Delta_{Student}$ 16.07% 15.28% 27.77% 21.31% 26.71% 19.41% 20.03% 16.55% 8.92% 6.38% 14.19% 8.23%
Weather Student 0.234 0.294 0.232 0.289 0.237 0.291 0.221 0.278 0.235 0.293 0.239 0.278
+TimeDistill 0.220 0.270 0.222 0.267 0.225 0.271 0.220 0.267 0.221 0.272 0.222 0.266
$\Delta_{Student}$ 5.83% 8.02% 4.53% 7.53% 5.22% 7.04% 0.53% 4.10% 5.95% 7.17% 7.19% 4.09%

### v2

@@ 6. Conclusion and Future work
We propose TimeDistill, a cross-architecture KD framework enabling lightweight MLP to surpass complex teachers. By distilling multi-scale and multi-period patterns, TimeDistill efficiently transfers temporal and frequency-domain knowledge to student MLP. Theoretical interpretations and comprehensive experiments confirm its versatility and effectiveness. Future work includes distilling from advanced time series models, e.g. time series foundation models, and incorporating multivariate patterns.
Appendix A Theoretical Understanding from Data Augmentation Perspective
[…]
@@ Appendix A Theoretical Understanding from Data Augmentation Perspective
(13) $\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug}$
when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (mixup) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@
where $x_{0}=Y$ and $x_{m+1}=\hat{Y}_{t}^{(m)}$ for $m=0,\dots,M$. By Jensen’s inequality (jensen1906fonctions), for a convex function $f$, we have:
@@
(14) $\mathcal{L}_{sup}+\eta\mathcal{L}_{period}\geq\mathcal{L}_{aug}$
@@
When the temperature $\tau$ goes to $\infty$, minimizing KL divergence between two distributions is equivalent to minimizing the MSE error (kim2021comparing). Thus, we have
@@
and the log sum inequality (thomas2006elements) states that
@@
Table 6. Dataset characteristics for time series long-term forecasting task.
@@
*The forecastability is borrowed from TimeMixer (timemixer). A larger value indicates better predictability.
Appendix B Implementation Details
Datasets Details. We conduct experiments to evaluate the performance and efficiency of TimeDistill on eight widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following the pipeline in (itransformer; timemixer; moderntcn). Detailed descriptions of these datasets are provided in Table 6.
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following (itransformer; moderntcn; timemixer; micn; timesnet; autoformer; informer; dlinear):
(15) $\text{MSE}=\frac{1}{S\times C}\sum_{i=1}^{S}\sum_{j=1}^{C}(Y_{ij}-\hat{Y}_{ij})^{2},$
(16) $\text{MAE}=\frac{1}{S\times C}\sum_{i=1}^{S}\sum_{j=1}^{C}|Y_{ij}-\hat{Y}_{ij}|.$
@@ Appendix B Implementation Details
Table 7. Experiment configuration of TimeDistill.
@@
Experiment Details. All experiments are implemented in PyTorch (paszke2019pytorch) and conducted on a cluster equipped with NVIDIA A100, V100, and K80 GPUs. The teacher models are trained using their default configurations as reported in their respective papers. For the student MLP model, we employ a combination of a decomposition scheme (dlinear; autoformer; fedformer) and a two-layer MLP with hidden dimension $D$ as the architecture. When using TimeDistill for distillation, the teacher model is frozen, and only the student MLP is trained. The initial learning rate is set to 0.01, and the ADAM optimizer (kingma2014adam) is used with MSE loss for the training of the student model. We apply early stopping with a patience value of 5 epochs. The batch size is set to 32. The temperature for multi-period distillation is set to $\tau=0.5$. The number of scales is set to $M=3$. We use two types of normalization methods, i.e. non-stationary (non-stationary) and revin (revin). We perform a hyperparameter search for $\alpha$ and $\beta$ within the range {0.1, 0.5, 1, 2}. We conduct hyperparameter sensitivity analysis in Appendix H. For consistency, TimeDistill defaults to using ModernTCN as the teacher and fixes the look-back window length to 720. Additional experiments with other teacher models and different look-back window lengths are presented in Section 5.3. Additional detailed model configuration information is presented in Table 7. We also report the standard deviation of TimeDistill performance under five runs with different random seeds in Table 8, which exhibits that the performance of TimeDistill is stable.
[…]
Table 8. Robustness of TimeDistill performance. The results are obtained from five random seeds, averaged over four prediction lengths.
Dataset ECL ETTh1 ETTh2 ETTm1 ETTm2 Solar Traffic Weather
[…]
Appendix C Ensuring Fair Comparison
To ensure a fair comparison, we have taken the following measures. First, we treat normalization (either non-stationary normalization or REVIN) as a hyperparameter and follow the original implementation of each baseline to obtain its optimal performance. Second, we reproduce all baselines using the widely adopted GitHub repository “Time-Series-Library” and implement TimeDistill within the same framework to provide a unified and fair evaluation environment. Third, we set drop_last=False so that all methods are evaluated on the exact same test set. Finally, the only difference between the baseline MLP and TimeDistill is the inclusion of the distillation loss, which confirms that the observed performance improvements come solely from the distillation process.
Appendix D Further Illustration on Motivation
TimeDistill does not rely on the assumption that the student model (MLP) must be weaker than the teacher model. While MLPs may sometimes match or even outperform teacher models, our method does not assume that MLPs are generally weaker. Instead, knowledge distillation is effective for two main reasons:
Empirical complementary strengths: As outlined in Section 3, our analysis shows that MLPs benefit from distilling complementary knowledge from teacher models. As shown in Figure 5, the win ratio is high, which indicates that teacher and student models have different strengths on different subsets of data. Importantly, the teacher model does not need to outperform the MLP; it is sufficient for the teacher to provide additional or diverse information that the MLP can exploit to improve its own performance.
Theoretical justification: We provide a theoretical interpretation of TimeDistill in Section 4.3, where we view it through a data augmentation perspective. This mixup-based data augmentation interpretation explains the generalization benefits of distillation, independent of the performance gap between teacher and student models.
Appendix E Implementation Details for Preliminary Study.
@@ Appendix D Further Illustration on Motivation
(17) $\mathbf{\hat{Y}}^{m}=\mathrm{Conv}(\mathbf{\hat{Y}}^{m-1},\mathrm{stride}=2),$
@@
Table 9. Performance promotion obtained by our TimeDistill framework with different teacher models. $\Delta_{MLP}$ ($\Delta_{Teacher}$) represents the performance promotion between MLP+TimeDistill and a trained MLP (Teacher). We report the average performance from all prediction lengths.
@@
((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@
Appendix F Different Teachers Analysis
@@
On Traffic dataset, which involves a large number of channels (861 channels) and is particularly challenging for channel-independent models like the student MLP, TimeDistill achieves the second-best performance, improving MSE by 10.30% when the teacher is iTransformer. This is likely due to TimeDistill can implicitly distill iTransformer’s knowledge of modeling multivariate correlations to student MLP. We further explore whether TimeDistill can implicitly learn multi-variate correlation via distillation in Appendix 9.
@@
Appendix G Different Students Analysis
@@
Appendix H Hyperparameter Sensitivity Analysis
@@
Figure 10. Sensitivity Analysis Result of Predcition Level $\alpha$ and Feature Level $\beta$ Distillation on Different Datasets.
@@
Figure 11. Hyperparameter sensitivity with respect to the number of scales $M$ and temperature $\tau$ on ECL dataset. The results are recorded with the lookback window length T = 720 and averaged across all prediction window lengths $S~\in\{96,192,336,720\}$.
Table 10. Results of directly match the patterns in ground truth time series on eight datasets. The look-back length is set to be consistent at 720 for fair comparison. The MSE and MAE metrics are averaged from all prediction lengths $S\in\{96,192,336,720\}$.
@@
Table 11. Static and runtime metrics of TimeDistill, MLP, and other mainstream models on the ECL dataset. For fair comparison, the look-back length for each model is set to be consistent at 720 and the batch size is set to be 16. The static and runtime metrics are averaged from all prediction lengths $S~\in\{96,192,336,720\}$.
@@
iTransformer (itransformer) 0.163 0.258 3.6377 5,276,496
ModernTCN (moderntcn) 0.167 0.262 6.2233 132,075,892
TimeMixer (timemixer) 0.165 0.259 7.7662 5,064,993
PatchTST (patchtst) 0.165 0.266 2.9802 24,949,584
MICN (micn) 0.181 0.293 2.715 60,215,726
Crossformer (crossformer) 0.203 0.299 8.7768 2,264,036
Fedformer (fedformer) 0.274 0.376 32.2642 16,827,399
Autoformer (autoformer) 0.238 0.347 196.5266 14,914,398
@@
Appendix I Directly Match the Patterns in Ground Truth Time Series
@@
(18) $\mathcal{L}^{\prime}_{sup}=\mathcal{L}_{scale}^{sup}+\mathcal{L}_{period}^{sup}.$
@@
Appendix J Efficiency Analysis
[…]